carlos and maria drove a total of 280 miles in 5.5 hours. carlos drove the forst part of the trip and averaged 49 miles per hour. maria drove the remainder of the trip and avereged 54 miles per hour. for approximately how many hours did maria drive? round your answer to the nearset tenth if necessary.
Found 2 solutions by nerdybill, checkley79:
Answer by nerdybill(7384)

carlos and maria drove a total of 280 miles in 5.5 hours. carlos drove the forst part of the trip and averaged 49 miles per hour. maria drove the remainder of the trip and avereged 54 miles per hour. for approximately how many hours did maria drive? round your answer to the nearset tenth if necessary.
.
Apply the "distance formula" of d=rt
.
Let h = number of hours Maria drove
then
5.5-h = number of hours Carlos drove
.
49(5.5-h) + 54h = 280
269.5-49h + 54h = 280
269.5 + 5h = 280
5h = 10.5
h = 2.1 hours

Answer by checkley79(3341)

D=RT
280=49X+53(5.5-X)
280=49X+291.5-53X
280-291.5=-4X
-4X=-11.5
X=-11.5/4
X=2.9 HOURS FOR THE 49 MPH. TRIP.
5.5-2.9=2.6 HOURS FOR MARIA'S DRIVE.